# Incentivize Contribution and Learn Parameters Too: Federated Learning with Strategic Data Owners

### v2

@@ 9             Send $s_{i}^{t+1},d_{i}^{t+1}$ to the center
Remark. Note that the choice of step sizes $\gamma$ and $\nu$ scale inversely with $m$ and $n$ respectively. Moreover, we impose certain restrictions on $\lambda$ and $\tilde{\lambda}$, which may be restrictive. In Section 4, we remove such issues and propose a learning algorithm with choices of $\gamma$ and $\nu$ that do not scale with $m$ or $n$.
Remark. We emphasize that in experiments, sufficiently small and constant values of $\gamma$ and $\nu$ work. Hence, the implicit choice obtained in the above theorem is an artifact of our theoretical analysis, and is not a reflection of practical scenarios.
In Section A.3, we provide an example of an application of this setting where the NE is different from the social optima. Therefore, it is imperative to ask for a method where the social optima can be obtained in a NE. We address this in the following section.
4 Welfare maximization with monetary transfers
@@
In this section, we consider the scenario where the center can make a payment $p_{i}(s_{i},s_{-i})$ to agent $i$ to alter her utility function. Therefore, the effective cost of agent $i$ becomes $z_{i}(s_{i},s_{-i})=c_{i}(s_{i})-p_{i}(s_{i},s_{-i})$. Note that, the payment can be either positive or negative, which determines whether the agent is subsidized or taxed respectively. It is reasonable to expect that the data contributors of the federated learning process are subsidized for their data contribution while the data consumers are charged payment for obtaining the learned parameter. We assume that the derivatives of the accuracy $a_{i}$ and cost function $c_{i}$ are bounded. We choose the following payment function:
$p_{i}(s)=\beta(s_{i}-\frac{\displaystyle 1}{\displaystyle n-1}\sum_{j\neq i}s_{j}),$ (7)
@@
$u_{i}(w,s)=a_{i}(w,s)-c_{i}(s_{i})+\beta(s_{i}-\frac{\displaystyle 1}{\displaystyle n-1}\sum_{j\neq i}s_{j}).$ (8)
@@ 4.1 Convergence to the NE $(w^{\textnormal{OPT}},s^{\max})$
In this setup, we first prove that the utility of every agent $i$ can be made strictly increasing in her own contribution $s_{i}$ in the following manner.
@@
The derivative of $a_{i}(w,s)$ with respect to $s_{i}$ is bounded away from $\infty$- ∞, i.e., $\frac{\partial}{\partial s_{i}}a_{i}(w,s)\geqslant-\tau$ for some $0\leqslant\tau<\infty$, for all $i\in N,w$, and $s$. Moreover, the cost functions $c_{i}(.)$ has bounded derivatives, i.e., $c^{\prime}_{i}(s_{i})\leqslant\zeta$, for all $s_{i}$ and $i$.
Note that the above assumptions are pretty mild. Previous works assume even stronger assumptions (like convexity) on $a_{i}$ and $c_{i}$ (see [25]). We show that the above bounds on the derivatives of $a_{i}$ and $c_{i}$ are sufficient to claim our result.
@@
Suppose Assumption 2 holds and we use the payment function given by Equation 7. Then, the utility of every agent is increasing in her own data contribution, i.e., $\frac{\partial}{\partial s_{i}}u_{i}(w,s_{i},s_{-i})>0,\forall s_{i}\in(0,s_{i}^{\max}),i\in N,\text{ and }w$ provided $\beta>\zeta+\tau$.
Remark (Knowledge of $\beta$). The knowledge of $\beta$ is required for theoretical tractability only. In the experiments, we do not assume any knowledge of $\beta$, rather we tune this hyperparameter. We observe that uniformly across all datasets, if $\beta$ is chosen larger than a threshold, we obtain the best performance of our proposed algorithms which also validates the above theoretical requirement.
This lemma implies that even in Algorithm 1 if we apply the above payment, $s$ will converge to the maximum value $s^{\max}$. However, unlike Algorithm 1, in this section we provide an algorithm which gives the step sizes of the gradient ascents in a more concrete manner and is, therefore, a superior one. In order to keep the convergence rate same as Theorem 1, we assume that the negative [^2] We negate the welfare so that we can consider functions as convex to apply the results of convex analysis easily. social welfare function at $s^{\max}$, given by $f(w,s^{\max})=-\frac{1}{n}\sum_{i\in N}a_{i}(w,s^{\max})$, is $M$-smooth and $\nu$-strictly convex in $w$. These properties are formally defined below.
@@ 11       Center broadcasts $w^{t}$
13             Compute local gradient: $d_{i}^{t+1}=\nabla_{w}\>a_{i}(w^{t},s_{i}^{\max},s_{-i}^{\max})$
@@ 14             Send $d_{i}^{t+1}$ to the center
We propose a two-phase algorithm given by Algorithm 2. The algorithm, in the first phase, incentivizes the agents to contribute $s^{\max}$, and in the second phase, converges to $w^{\texttt{{OPT}}}$. Note that in the first phase of Algorithm 2, every agent runs a gradient ascent step. Thanks to the increasing utility property (Lemma 1), we have an increasing sequence of $s_{i}^{t}$ for all $i\in N$. Moreover, from the definition of $[g(w^{0},s^{t},\mu^{t})]_{i}$ it is ensured that $s_{i}^{\max}$ is the fixed point of this gradient ascent update. Hence after a finite number of iterations, every agent reaches $s_{i}^{\max}$. Note that in this phase the model parameter $w^{0}$ remains unchanged.
@@
We now provide the convergence guarantees of Algorithm 2. Before that, let us discuss the assumptions as discussed in Section A.3.
@@
The smoothness and strong convexity assumptions have featured in several previous papers on FL [16, 37]. The main result of this section is as follows.
@@
Suppose Assumptions  2 and 3 hold and we consider the utility function with payment scheme given in Equation 8 with $\beta>\zeta+\tau$. Then, for every $w\in\mathbb{R}^{m}$ and $s\in\prod_{i\in N}S_{i}$, 2P-UPBReD (Algorithm 2) converges to the Nash equilibrium $(w^{\texttt{{OPT}}},s^{\max})$ and is budget balanced. In particular, for any given $\epsilon>0$, we have
$\frac{1}{n}\sum_{i\in N}a_{i}(w^{\texttt{{OPT}}},s^{\max})-\frac{1}{n}\sum_{i\in N}a_{i}(w^{T},s^{\max})<\epsilon$

### v3

@@ 9      Send $s_{i}^{t+1},d_{i}^{t+1}$ to the center
Remarks. The step sizes $\gamma$ and $\nu$ in our analysis scale inversely with $m$ and $n$, and certain restrictions on $\lambda$ and $\tilde{\lambda}$ may appear restrictive. In Section 4, we address these issues and propose a learning algorithm with step sizes independent of $m$ and $n$.
In practice, however, we observe that sufficiently small constant values of $\gamma$ and $\nu$ work well, indicating that the scaling requirement in the theorem is merely a theoretical artifact rather than a practical limitation.
While agents strategically choose $s_{i}$ to maximize their utility, we assume truthful reporting of these choices. In UPBReD, agents report $s_{i}^{t}$ and $d_{i}^{t}$ in each round. The next subsection introduces payments to ensure truthful reporting of $s_{i}^{t}$ and $d_{i}^{t}$.
3.1 Truthful elicitation with payment rules
So far, we have assumed that the agents report their type truthfully in every round, i.e. if an agent is training on $s_{i}^{t}$ amount of data at time $t$ the agent reports $\theta_{i}=(s_{i}^{t},d_{i}^{t})$ truthfully. However, an agent can misreport their strategy and corresponding gradient update to improve their utility as observed in the example below.
Example 2.
We perform some experiments on the CIFAR data set to show that agents have an incentive to misreport. Consider $2$ agents each having a set of $480$ data points. $u_{i}=56-L(w,s)-c_{i}(s_{i})$, where $L(w,s)$ is the cross entropy loss computed on the dataset. The costs for the agents are $c_{1}(s_{1})=0.2s_{i},c_{2}(s_{2})=0.1s_{2}$. When agents truthfully report $s_{i},d_{i}$ in UPBReD they observe a social welfare of $15.5$, with $s_{1}*=270,s_{2}*=165$ this yields a utility $u_{1}=510$. When agent misreports $s_{1}^{t}=240$ and reports $d_{1}^{t}=0_{m}$ in every training round yielding a utility of $u_{i}=64$. Clearly agent $1$ is better off by misreporting their type. ∎
A well established goal in social choice theory is to choose the allocation that maximizes the social welfare (Green and Laffont, 1977). In our setting this maximizing allocation for a given type profile $\theta^{t}$ at round $t$ will be
$a_{*}({\theta}^{t})=\operatorname*{\arg\!\max}_{w\in\mathbb{R}^{m}}{\sum_{i\in N}v_{i}(w,s^{t}})$ (12)
However, in our federated learning setting, we cannot compute this welfare maximizing allocation. Instead, we use $a_{w}^{t}$, which iteratively takes gradient ascent steps towards the welfare maximizing allocation as shown in Algorithm˜1. The initial allocation is chosen as $w^{0}$. And future allocations are computed iteratively as $a_{w}^{t}(\theta^{t})=w^{t}(s^{t},d^{t})=w^{t-1}+\frac{\eta}{n}\sum_{i\in N}d_{i}^{t}$. Due to this difference, we need an additional assumption for truthfulness. We assume the difference between the welfare at our computed allocation and that of the maximum at truthful reports is bounded by a constant.
Assumption 2.
$\sum_{i\in N}v_{i}(a_{*}(\theta^{t}),s^{t})-\sum_{i\in N}v_{i}(a_{w}^{t}(\theta^{t}),s^{t})<\Gamma,\>\forall\theta^{t}\in\Theta,\ \forall t=1,\ldots,T$.
Consider the following payment rule:
$p_{i}^{*}(\hat{\theta}^{t},\hat{v}^{t})=\sum_{j\neq i}\hat{v}_{j}^{t}-\phi(\hat{v}_{i}^{t},v_{i}(a_{w}^{t}(\hat{\theta}^{t}),\hat{s}^{t}))-h_{i}(\hat{\theta}_{-i}^{t})$ (13)
where $\phi$ is given by,
$\textstyle\phi\left(\hat{v}^{t}_{i},v_{i}(a_{w}^{t}(\hat{\theta}^{t}),\hat{s}^{t})\right)$ $\textstyle=0\quad\text{if }\hat{v}^{t}_{i}=v_{i}(a_{w}^{t}(\hat{\theta}^{t}),\hat{s}^{t})$
$\textstyle=\Gamma+(\hat{v}^{t}_{i}-v_{i}(a_{w}^{t}(\hat{\theta}^{t}),\hat{s}^{t}))^{2}\quad\text{otherwise}$
T-UPBReD initializes a parameter $w^{0}$ and agents begin with an initial vector $s^{0}$. In every subsequent round T-UPBReD follows the following steps at round $t$.
  1. 1.
    The center computes and shares a parameter $w^{t-1}$ computed in the previous round with the agents.
  2. 2.
    Agents compute the gradients $d_{i}^{t}$ with respect to $w^{t-1}$ on their chosen dataset size $s_{i}^{t-1}$ then strategically pick their dataset size for the next round $s_{i}^{t}$ using Equation˜3, which forms the true type of the agent $\theta^{t}_{i}=(s_{i}^{t},d_{i}^{t})$.
  3. 3.
    Agents share their type as $\hat{\theta}_{i}^{t}=(\hat{s}_{i}^{t},\hat{d}_{i}^{t})$ which can be different from their true type. The center computes an allocation using the reported types using $a_{w}^{t}$ consistent with Equation˜4.
  4. 4.
[…]
  5. 5.
    Agents report their valuation as $\hat{v}_{i}^{t}$ which may be different from the true valuations. The center then computes payments $p_{i}^{*}(\hat{\theta}^{t},\hat{v}_{i}^{t})$ for each agent.
With the given allocation rule $a_{w}^{t}$ and payment rule $p$ described in Equation˜13 for a true type and valuation profile $\theta^{t},v^{t}$, a reported type and valuation profile $\hat{\theta}^{t},\hat{v}^{t}$, the utility realised by agent $i$ at round $t$ is given by $u_{i}(\hat{\theta}^{t},\hat{v}^{t}|\theta^{t},v^{t})=v_{i}(a_{w}^{t}(\hat{\theta}^{t}),s^{t})-z_{i}(s_{i}^{t},\hat{s}_{-i}^{t})+p_{i}^{*}(\hat{\theta}^{t},\hat{v}_{i}^{t})$. The algorithm is detailed in Algorithm˜2.
1
Output: $w^{T}$
2
3for $t=0$ to T1T-1 do
4 Center broadcasts $w^{t},s^{t}$
5
6 foreach agent $i\in N$ in parallel do
7    $s^{t+1}_{i}=s^{t}_{i}+\gamma[g(w^{t},s^{t},\mu^{t})]_{i}$
8      Compute local gradient: $d_{i}^{t+1}=\nabla_{w}\>v_{i}(w^{t},s_{i}^{t},s_{-i}^{t})$
9      Send $\tilde{s}_{i}^{t+1},\tilde{d}_{i}^{t+1}$ to the center
10
11
12 Center updates: $w^{t+1}=w^{t}+\frac{\eta}{n}\sum_{i\in N}\tilde{d}_{i}^{t+1}$
13
14 Center broadcasts $w^{t+1}$
15 foreach agent $i\in N$ in parallel do
16      Observes $v_{i}(w^{t+1},\tilde{s}^{t+1})$
17     Send $\hat{v}_{i}^{t}=\hat{a}_{i}$
18
19 Center computes payments using Equation˜13 at $w^{t}$
20
21return $w^{T}$
22
Algorithm 2 Truthful Updated Parameter Best Response Dynamics (T-UPBReD)
With this algorithm, we can show the following result.
Theorem 2.
Under ˜2, Algorithm˜2 is EPIC in every round.
Proof.
Consider the utility of agent $i$ at time $t$, whose true type and valuation are $\theta_{i}^{t}=(s_{i}^{t},d_{i}^{t}),v_{i}^{t}$ respectively, misreporting their type and valuation, $\hat{\theta}^{t}_{-i}=(\hat{s}_{i}^{t},\hat{d}_{i}^{t})$, $\hat{v}_{i}^{t}$ while other agents are truthful $\theta_{-i^{t}}=(s_{-i}^{t},d_{-i}^{t})$.
$u_{i}(\hat{\theta}_{i}^{t},\theta_{-i}^{t},\hat{v}_{i}^{t},v_{-i}^{t}|\theta^{t},v^{t})$
$=v_{i}(a_{w}^{t}(\hat{\theta}_{i}^{t},\theta_{-i}^{t}),s^{t})-z_{i}(s_{i}^{t},s_{-i}^{t})+p_{i}(\hat{\theta}^{t}_{i},\theta_{-i}^{t},\hat{v}_{i}^{t},v_{-i}^{t})$
Expanding the terms as defined in Equation˜13
$u_{i}(\hat{\theta}_{i}^{t},\theta_{-i}^{t},\hat{v}_{i}^{t},v_{-i}^{t}|\theta^{t},v^{t})$
$=v_{i}(a_{w}^{t}(\hat{\theta}_{i}^{t},\theta_{-i}^{t}),s^{t})-z_{i}(s_{i}^{t},s_{-i}^{t})+\sum_{j\neq i}v_{j}(a_{w}^{t}(\hat{\theta}_{i}^{t},\theta_{-i}^{t}),s^{t})$
$-\Gamma-(\hat{v}^{t}_{i}-v_{i}(a_{w}^{t}(\hat{\theta}^{t}),\hat{\theta}^{t}))^{2}-h_{i}(\theta_{-i}^{t})$
Since $(\hat{v}^{t}_{i}-v_{i}(a_{w}^{t}(\hat{\theta}^{t}),\hat{\theta}^{t}))^{2}$ is always non-positive, we get
$u_{i}(\hat{\theta}_{i}^{t},\theta_{-i}^{t},\hat{v}_{i}^{t},v_{-i}^{t}|\theta^{t},v^{t})$
$\leq\sum_{j\in N}v_{j}(a_{w}^{t}(\hat{\theta}^{t}_{i},\theta_{-i}^{t}),s^{t})-z_{i}(s_{i}^{t},s_{-i}^{t})-\Gamma-h_{i}(\theta_{-i}^{t})$
As social welfare is maximized by Equation˜12
$u_{i}(\hat{\theta}_{i}^{t},\theta_{-i}^{t},\hat{v}_{i}^{t},v_{-i}^{t}|\theta^{t},v^{t})$
$\leqslant\sum_{j\in N}v_{j}(a_{*}(\theta^{t}),s^{t})-z_{i}(s_{i}^{t},s_{-i}^{t})-\Gamma-h_{i}(\theta_{-i}^{t})$
Using ˜2
$u_{i}(\hat{\theta}_{i}^{t},\theta_{-i}^{t},\hat{v}_{i}^{t},v_{-i}^{t}|\theta^{t},v^{t})$
$\leqslant\sum_{j\in N}v_{j}(a_{w}^{t}(\theta^{t}),s^{t})+\Gamma-\Gamma-z_{i}(s_{i}^{t},s_{-i}^{t})-h_{i}(\theta_{-i}^{t})$
$=\sum_{j\in N}v_{j}(a_{w}^{t}(\theta^{t}),s^{t})-z_{i}(s_{i}^{t},s_{-i}^{t})-h_{i}(\theta_{-i}^{t})$
$=u_{i}(\theta^{t},v^{t}|\theta^{t},v^{t})$
3.2 Application of Theorem˜1: effective costs are only personal
One special but practical case of the above setup is when the effective cost is only borne by the agent. Mathematically, it is represented as $z_{i}(s_{i},s_{-i})=c_{i}(s_{i}),\forall s_{i}\in S_{i},s_{-i}\in S_{-i},\forall i\in N$. We assume convex cost functions $c_{i},\forall i\in N$, which is a commonly used assumption Li and Raghunathan (2014) in the literature. This setup is consistent with the assumptions of ˜1, and according to Theorem˜1, Algorithm˜1 converges to a Nash equilibrium, which we denote as $(w^{*},s^{*})$. However, it is possible that none of the following things happen: (i) $s^{*}\neq s^{\max}$, i.e., the agents do not contribute their entire data for the federated learning process, leading to a suboptimal learning, or (ii) $w^{*}\neq w^{\texttt{{OPT}}}$ (see Equation˜2 for the definition), which is neither the objective of the center nor the agents. The following example shows such an instance.
Example 3 (NE different from socially optimal).
Consider the federated learning setup with two agents that have linear costs and are trying to learn a model with two parameters. The valuation function is identical and is given by $v_{i}(w,s)=1-\frac{(1-w_{1})^{2}+(2-w_{2})^{2}}{s_{1}+s_{2}},i=1,2$. The cost functions for agents $1$ and $2$ are given by $c_{1}(s_{1})=0.04\cdot s_{1}$ and $c_{2}(s_{2})=0.02\cdot s_{2}$ respectively. The agents’ strategy sets are $S_{i}=[0,5],i=1,2$. Notice that, at $s=s^{\max}=(5,5)$, the value of $w^{\texttt{{OPT}}}$ (see Equation˜2) is $(1,2)$. It yields an optimal social welfare of $2$. However, this is not an NE, since the derivatives $\partial u_{i}/\partial s_{i}|_{s_{i}^{\max}}<0$ for both $i=1,2$. Hence, the best response of each agent is to reduce $s_{i}$. However, from this point, with the choice of $\gamma=0.25$ and $\eta=0.25$, Algorithm˜1 converges to an NE profile of $w^{*}=(0.5,1.5),s^{*}=(0,5)$ that yields a social welfare of $1.8$. ∎
4 Welfare maximization at Nash equilibrium
@@
In this section, we consider the scenario where the center can make a payment $p_{i}(s_{i},s_{-i})$ to agent $i$ to alter its utility function. Therefore, the effective cost of agent $i$ becomes $z_{i}(s_{i},s_{-i})=c_{i}(s_{i})-p_{i}(s_{i},s_{-i})$. Note that, the payment can be either positive or negative, which determines whether the agent is subsidized or taxed respectively. It is reasonable to expect that the data contributors of the federated learning process are subsidized for their data contribution while the data consumers are charged payment for obtaining the learned parameter. We assume that the derivatives of the valuation $v_{i}$ and cost function $c_{i}$ are bounded. We choose the following payment function:
$p_{i}(s)=\beta\left(s_{i}-\frac{\textstyle 1}{\textstyle n-1}\sum_{j\neq i}s_{j}\right),$ (14)
@@
$u_{i}(w,s)=v_{i}(w,s)-c_{i}(s_{i})+\beta(s_{i}-\frac{\textstyle 1}{\textstyle n-1}\sum_{j\neq i}s_{j}).$ (15)
@@ 4.1 Convergence to the NE $(w^{\textnormal{OPT}},s^{\max})$
In this setup, we first prove that the utility of every agent $i$ can be made strictly increasing in their own contribution $s_{i}$ in the following manner.
@@
The derivative of $v_{i}(w,s)$ with respect to $s_{i}$ is bounded away from -$\infty$, i.e., $\frac{\partial}{\partial s_{i}}v_{i}(w,s)\geqslant-\tau$ for some $0\leqslant\tau<\infty$, for all $i\in N,w$, and $s$. Moreover, the cost functions $c_{i}(.)$ has bounded derivatives, i.e., $c^{\prime}_{i}(s_{i})\leqslant\zeta$, for all $s_{i}$ and $i$.
Note that the above assumptions are pretty mild. Previous works assume even stronger assumptions (like convexity) on $v_{i}$ and $c_{i}$ (see (Murhekar et al., 2023)). We show that the above bounds on the derivatives of $v_{i}$ and $c_{i}$ are sufficient to claim our result.
@@
Suppose Assumption 3 holds and we use the payment function given by Equation˜14. Then, the utility of every agent is increasing in its own data contribution, i.e., $\frac{\partial}{\partial s_{i}}u_{i}(w,s_{i},s_{-i})>0,\forall s_{i}\in(0,s_{i}^{\max}),\forall i\in N,\text{ and }\forall w$ provided $\beta>\zeta+\tau$.
Proof.
From Equation˜15, we get the derivative of the utility function for $s_{i}\in(0,s_{i}^{\max})$ (note that this holds only for the interior of $S_{i}$, the derivative at the boundaries are zero by definition of function $g$ in Section˜3) to be $g(w,s,\mu)]_{i}=\frac{\partial}{\partial s_{i}}u_{i}(w,s)=\frac{\partial}{\partial s_{i}}v_{i}(w,s)-c_{i}^{\prime}(s_{i})+\beta$, where $c_{i}^{\prime}(s_{i})=\frac{d}{ds_{i}}c_{i}(s_{i})$. Using Assumption 3 and choosing in Equation˜14 the parameter $\beta>\zeta+\tau$, we have
$[g(w,s,\mu)]_{i}$ $=\frac{\partial}{\partial s_{i}}v_{i}(w,s)-c_{i}^{\prime}(s_{i})+\beta\geqslant-\tau-\zeta+\beta>0.$
The last inequality holds since $\beta>\zeta+\tau$, $\forall i\in N$. ∎
Remark (Knowledge of $\beta$). The knowledge of $\beta$ is required for theoretical tractability only. In the experiments, we do not assume any knowledge of $\beta$, rather we tune this hyperparameter. We observe that uniformly across all datasets, if $\beta$ is chosen larger than a threshold, we obtain the best performance of our proposed algorithms, which also validates the above theoretical requirement.
This lemma implies that even in Algorithm˜1 if we apply the above payment, $s$ will converge to the maximum value $s^{\max}$. However, unlike Algorithm˜1, in this section we provide an algorithm which gives the step sizes of the gradient ascents in a more concrete manner and is, therefore, a superior one. In order to keep the convergence rate same as Theorem˜1, we assume that the negative [^2] We negate the welfare so that we can consider functions as convex to apply the results of convex analysis easily. social welfare function at $s^{\max}$, given by $f(w,s^{\max})=-\frac{1}{n}\sum_{i\in N}v_{i}(w,s^{\max})$, is $M$-smooth and $\nu$-strictly convex in $w$. These properties are formally defined below.
@@ 11 Center broadcasts $w^{t}$
13      Compute local gradient: $d_{i}^{t+1}=\nabla_{w}\>v_{i}(w^{t},s_{i}^{\max},s_{-i}^{\max})$
@@ 14      Send $d_{i}^{t+1}$ to the center
We propose a two-phase algorithm given by Algorithm˜3. The algorithm, in the first phase, incentivizes the agents to contribute $s^{\max}$, and in the second phase, converges to $w^{\texttt{{OPT}}}$. Note that in the first phase of Algorithm˜3, every agent runs a gradient ascent step. Thanks to the increasing utility property (Lemma 1), we have an increasing sequence of $s_{i}^{t}$ for all $i\in N$. Moreover, from the definition of $g(w^{0},s^{t},\mu^{t})]_{i}$ it is ensured that $s_{i}^{\max}$ is the fixed point of this gradient ascent update. Hence after a finite number of iterations, every agent reaches $s_{i}^{\max}$. Note that in this phase the model parameter $w^{0}$ remains unchanged.
@@
We now provide the convergence guarantees of Algorithm˜3. Before that, let us discuss the necessary assumptions.
@@
The smoothness and strong convexity assumptions have featured in several previous papers on FL (Karimireddy et al., 2020b; Yin et al., 2018). The main result of this section is as follows.
@@
Suppose Assumptions  3 and 4 hold and we consider the utility function with payment scheme given in Equation˜15 with $\beta>\zeta+\tau$. Then, for every $w^{0}\in\mathbb{R}^{m}$ and $s^{0}\in\prod_{i\in N}S_{i}$, 2P-UPBReD (Algorithm˜3) converges to the Nash equilibrium $(w^{\texttt{{OPT}}},s^{\max})$ and is budget balanced. In particular, for any given $\epsilon>0$, we have
$\frac{1}{n}\sum_{i\in N}v_{i}(w^{\texttt{{OPT}}},s^{\max})-\frac{1}{n}\sum_{i\in N}v_{i}(w^{T},s^{\max})<\epsilon$